OpenStudy (asapbleh):

Find the component form of the vector v in 2-space that has the stated length and makes the stated angle theta with the positive x-axis.

OpenStudy (asapbleh):

\[\left| \left| v \right| \right|=3; \theta=\pi/4\]

OpenStudy (asapbleh):

so i used trig and its sqrt2/2,sqrt2/2. then do u just multiple the magnitude by the vector?